OpenStudy (anonymous):

so very very confusedled Write a system of two equations using x and y to represent the following problem. A chemist is mixing two solutions together. The chemist wants 100 gallons when complete. He is mixing a 24% solution with a 50% solution to get a 36% solution. How many gallons of each will he use? please please explain

OpenStudy (jennychan12):

oi. percents. x+y = 100 gallons 0.5x+0.36y = 1 entire solution solve for x and y

OpenStudy (anonymous):

so i need the two equations?

OpenStudy (jennychan12):

yeah lol it's called system of equations for a reason haa

OpenStudy (anonymous):

ok sooo what do i do now?

OpenStudy (jennychan12):

oh wait. sorry i told you wrong. i hate these types of problems. sorry i'm not sure what to do...

OpenStudy (jennychan12):

http://www.cliffsnotes.com/study_guide/Mixture-Problems.topicArticleId-254915,articleId-254911.html maybe this will help? look at example 2.

OpenStudy (jennychan12):

So you have 0.5x + 100 +0.24(100-x) = 0.36 <-- this is the hard part, setting up the equation Just solve for x. See attached table for a clearer "picture" of the situation.